OpenStudy (anonymous):

The hypotenuse of a 30-60-90 triangle has length 19. What is the length of the side opposite the 60 angle? If necessary, round your answer to two decimal places.

OpenStudy (anonymous):

http://openstudy.com/users/terenzreignz

terenzreignz (terenzreignz):

You know, that's not how you tag :D

OpenStudy (anonymous):

i know lol

OpenStudy (raden):

let the length of the side opposite the 60 angle is x. x/19 = sqrt(3)/2 solve for x = ...

OpenStudy (raden):

hint : multiply by 19 to both sides